OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-5, 10) and (-4,3)

OpenStudy (anonymous):

slope is \[\frac{y_2-y_1}{x_2-x_1}=\frac{3-10}{-4+5}=-7\]

OpenStudy (anonymous):

then use \[y-y_1=m(x-x_1)\] with \[m=-7,x_1=-5,y_1=10\] so get \[y-10=-7(x+5)\] \[y-10=-7x-35\] \[y=-7x-25\]